OpenStudy (anonymous):

For 9 years, Christina played goalie on a soccer team. Find the range, mode, and median of her saves per year: 23, 39, 68, 17, 39, 39, 25, 17, 45

OpenStudy (anonymous):

arrange the data first

OpenStudy (anonymous):

and then u will get median

OpenStudy (anonymous):

i need to range and mode and median

OpenStudy (anonymous):

@suzi20

OpenStudy (anonymous):

remember the mode is the most apparent number and the median is the middle value write it order

OpenStudy (anonymous):

From what I remember from least to greatest so 17, 17 ,23,25, 39 ,39 ,39 , 45 there now try to find the mode and median the range I cant help you with i was never really good at it.

OpenStudy (anonymous):

oops forgot 68 heh 68 is last kay.

OpenStudy (anonymous):

Hope this helps !!!!

OpenStudy (anonymous):

17, 17 ,23, 25, 39, 39 ,39 , 45, 68 Median = 39 (the data at the center after arranged from least to greatest) Mode = 39 (too), (mode is the most frequent appearing data) Mean=(17+17+23+25+39+39+39+45+68)/9

OpenStudy (anonymous):

exactly thank you imron i was just to lazy to type the answer ^-^

OpenStudy (anonymous):

What's the range? @imron07

OpenStudy (anonymous):

@sweetkit : I know, sometimes I feel the same way ;)

OpenStudy (anonymous):

xD oh @imron07 you shouldn't give her the answer it isn't doing any good wusu try to figure that out on you own kay i know you can do it!

OpenStudy (anonymous):

Sorry @owusulovesyou , @sweetkit doesn't allow me to give the answer xD Just what is the definition of range of set of data? (I count three ofs here)

OpenStudy (anonymous):

i said shouldn't never said you could!!! don't twist my word @imron07

OpenStudy (anonymous):

Heheh, sorry. Okay i'm leaving.

OpenStudy (anonymous):

.....kay sorry if i offended >///< oh well have an epic night and funtastic life!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

creepy ......

OpenStudy (anonymous):

wusu did you figure out the range yet?

OpenStudy (anonymous):

gah!!!! bye bye have fun with the range wusu!